OpenStudy (anonymous):

Given the points (2,1) and (3,4), answer the following. A) What is the slope of the line through these points? B) Write the equation in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1). C) Simplify your answer in part B and write the equation in slope-intercept form. Show all work.

OpenStudy (anonymous):

@goformit100

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

i got 3

hartnn (hartnn):

that is correct

OpenStudy (anonymous):

for part a

hartnn (hartnn):

slope point form y-y1=m(x-x1) where (x1,y1) is a point on the line

OpenStudy (anonymous):

so can u help me to get part b

hartnn (hartnn):

the formula was already given in question, just put values

OpenStudy (anonymous):

i need help with it

hartnn (hartnn):

lets take the point (2.1) y-1=3(x-2)

OpenStudy (anonymous):

i think its y = 3x

OpenStudy (anonymous):

and stuff added

hartnn (hartnn):

after u put your point, u get y-1=3(x-2) this is your 2nd part. now for 3rd part try to isolate y

OpenStudy (anonymous):

can u help me ?

OpenStudy (anonymous):

y=3x-5

hartnn (hartnn):

just add 1 to each side